one canned juice drink is 15% orange juice; another is 5% orange juice. How many liters of each should be mixed together in order to get 10L that is 7% orange juice?
Answer by mananth(16946)

percent ---------------- quantity
Juice I 15 ---------------- x liters
JuiceII 5 ---------------- 10 - x liters
Total 7.00% ---------------- 10

15x+5(10-x)=7*10

15x+50-5x=70
15x-5x=70-50
10x=20
/10
x=2 liters 15.00%
8liters 5.00%